# Between Circuits and Chomsky: Pre-pretraining on Formal Languages Imparts Linguistic Biases

### v1

@@ Abstract
Pretraining language models on formal languages can improve their acquisition of natural language, but it is unclear which features of the formal language impart an inductive bias that leads to effective transfer. Drawing on insights from linguistics and complexity theory, we hypothesize that effective transfer occurs when the formal language both captures dependency structures in natural language and remains within the computational limitations of the model architecture. Focusing on transformers, we find that formal languages with both these properties enable language models to achieve lower loss on natural language and better linguistic generalization compared to other languages. In fact, pre-pretraining, or training on formal-then-natural language, reduces loss more efficiently than the same amount of natural language. For a 1B-parameter language model trained on roughly 1.6B tokens of natural language, pre-pretraining achieves the same loss and better linguistic generalization with a 33% smaller token budget. We also give mechanistic evidence of cross-task transfer from formal to natural language: attention heads acquired during formal language pretraining remain crucial for the model’s performance on syntactic evaluations.
@@ 1 Introduction
Language models have achieved impressive performance on many tasks, but they remain data-hungry, requiring five to six orders of magnitude more data than humans to achieve human-level performance Warstadt et al. (2023); Gilkerson et al. (2017). This high data requirement presents challenges for training models in low-resource settings Zhong et al. (2024); Hettiarachchi et al. (2024), understanding how models can generalize given more human-like data constraints Wilcox et al. (2024), and improving models as original data becomes scarce Villalobos et al. (2022). Thus, an important frontier for language models is making training more data-efficient.
@@
A recently-explored approach for increasing data efficiency teaches models useful inductive biases by “pre-pretraining” them on formal languages before training on natural language Chiang and Lee (2021); McCoy and Griffiths (2023). In particular, Papadimitriou and Jurafsky (2023) show that within the Chomsky hierarchy, context-sensitive languages transfer best to natural language. However, results from circuit complexity show that transformers cannot learn next-token prediction for all context-sensitive languages, both in theory and practice Strobl et al. (2024); Merrill and Sabharwal (2023). Within all levels of the Chomsky hierarchy, some languages are harder for transformers to learn than others, if not impossible Merrill et al. (2023, 2024). These results call into question how positive transfer occurs, if a generalizing solution cannot always be learned. Which characteristics of formal languages make them effective for pre-pretraining?
In this work, we propose that optimal transfer from formal to natural language occurs at the intersection of two theoretical hierarchies: the Chomsky hierarchy of formal languages and the circuit complexity hierarchy that bounds transformer computational power (§3). Specifically, we hypothesize that effective pre-pretraining languages should be:
@@
    learnable by transformers in a length-generalizing way.
Our empirical results support this hypothesis (§4): pre-pretraining on languages with both the above properties outperforms pre-pretraining on either natural language or any of the other formal languages that we tested. Next, we show that when positive transfer occurs, the model reuses attention heads it learned during pre-pretraining, suggesting that mechanisms from pre-pretraining transfer to natural language (§5). We conclude by demonstrating our pre-pretraining method generalizes to 1B-parameter language model pretraining, increasing token efficiency by 33% (§6). In summary, our contributions are three-fold:
  1. 1.
    We propose a taxonomy of formal languages for pre-pretraining transformers (§3.2). Our framework reconciles constraints on formal languages from linguistic theory, as captured by the Chomsky hierarchy, and constraints from circuit complexity theory, as captured by the nascent circuit hierarchy Yang and Chiang (2024).
  2. 2.
    The best pre-pretraining languages we observe contain hierarchical structure and are learnable by transformers in a length-generalizing way (§4). We demonstrate that attention heads learned during pre-pretraining continue to be used after transfer to natural language (§5).
  3. 3.
    Finally, ablations reveal that hierarchical dependency structures are crucial for positive transfer. We also show that the benefit of pre-pretraining persists for 1B-parameter models (§6).
@@ 2.1 The Chomsky Hierarchy
The Chomsky hierarchy Chomsky (1959) is a nested classification of increasingly complex formal languages. This classification clarifies the kinds of computations needed to process formal structures resembling those found in human language. For example, regular languages, the least complex, can be recognized by finite-state automata; most phenomena in natural language phonology and morphology can be captured by regular languages. Regular languages are insufficient for syntax, however, as representing the hierarchical structure of natural language with a finite-state automaton would require infinitely many states Chomsky (1956). Subsequent works showed that compactly modeling natural-language syntax requires not only context-free but also context-sensitive grammars Shieber (1985). That said, only a select few phenomena in natural language require context sensitivity; those include cross-serial dependencies in Swiss German Bresnan et al. (1982) or anaphora; the rest can be modeled using context-free grammars Gazdar (1982).
@@
A classic context-free language is $k$-Dyck: the language of well-balanced parentheses with $k$ bracket types. For example, ([])[] is a valid $2$-Dyck string, where rounded and square parentheses are the two bracket types. $k$-Dyck is often taken as a canonical example of context-free hierarchical structure because any context-free language can be reduced to Dyck via one transformation (inverse homomorphism) and intersection with a regular language (Chomsky and Schützenberger, 1959).
@@
Removing the constraint that Dyck braces must be well-nested (but enforcing that every opening brace must be closed and vice versa) yields $k$-Shuffle Dyck, [^1] The name “Shuffle” comes from the fact that $k$-Shuffle Dyck can be defined by interleaving $k$ Dyck strings with different braces (Suzgun et al., 2019), as if by riffle shuffling. a minimal relaxation of $k$-Dyck that is strictly context-sensitive rather than context-free Suzgun et al. (2019); Strobl et al. (2024). Crossing braces in $k$-Shuffle Dyck can be thought of as a formal model of the cross-serial dependencies underlying aspects of language argued to be context-sensitive (Papadimitriou and Jurafsky, 2023).
@@ 2.2 The Circuit Hierarchy
Our focus in this work is on transformer language models. At each level of the Chomsky hierarchy, there are languages that a transformer cannot recognize (Merrill and Sabharwal, 2023; Liu et al., 2024; Strobl et al., 2024). Thus, the Chomsky hierarchy alone does not precisely capture how difficult a language is for transformers to learn: for instance, transformers can learn some context-free languages Butoi et al. (2025) and yet fail to learn other regular languages Merrill et al. (2024). To better understand the expressive power of transformers, recent work has analyzed formal languages within a circuit complexity hierarchy, native to what transformers can and cannot express Hao et al. (2022); Yang et al. (2024). Here, we will focus on two logics that emerge from this circuit complexity viewpoint: FO(M) Merrill and Sabharwal (2023) and C-RASP Yang and Chiang (2024).
@@
First-order logic with majority, or FO(M), is a provable upper bound on the languages that transformers can express: that is, any transformer can be converted into an FO(M) program that defines (or recognizes) the same language (Merrill and Sabharwal, 2023). FO(M) programs operate by computing counts over the number of indices in an input string that satisfy certain predicates. For example, $Q_{a}(i)$ is a basic predicate that checks whether input token $i$ is an $a$. The following FO(M) program uses $Q_{a}(i)$ to define the language of strings with exactly 3 $a$’s:
@@
Beyond this example, FO(M) can implement a rich variety of programs by nesting quantifiers and building complex predicates out of logical (,,$\wedge$,$\vee$,$\neg$ , ∨ , ¬) and arithmetic operators (+,=,<+,=,<+ , = , <). In particular, FO(M) can define the $k$-Dyck language for any $k\geq 1$ (Proposition A.6). For example, the following program defines $1$-Dyck:
@@
To finish the definition, we add a condition to enforce that any open and close brace paired by $\mathsf{depth}$ and $\mathsf{dindex}$ also match in their type (i.e., parenthesis or square brace). See Proposition A.7 for further details.
@@
While any transformer can be compiled into FO(M), it is not necessarily the case that any FO(M) program can be implemented by a transformer. C-RASP is a restriction of FO(M) designed to be a lower bound on what transformers can express: that is, if a language is definable in C-RASP, then there exists a transformer that recognizes it (Yang and Chiang, 2024). [^2] C-RASP is a fully well-defined variant of the Restricted Access Sequence Processing programming language (RASP; Weiss et al., 2021; Lindner et al., 2023). C-RASP works similarly to FO(M), but with a couple of restrictions. Most crucially, in C-RASP each predicate can only refer to one index variable $i$, whereas in FO(M), predicates can refer to two (or more) indices $i,j$ introduced by different quantifiers (for more detail, see Yang and Chiang, 2024). This means C-RASP can define (1) or (2) above, but not $k$-Dyck for $k\geq 2$, as C-RASP cannot express the $\mathsf{dindex}$ in (3).
Recent work has also suggested a connection between C-RASP and length generalization of transformers (Zhou et al., 2024; Huang et al., 2025): the definability of a language $L$ in C-RASP predicts whether transformers can reliably length-generalize when trained on strings from $L$. One interpretation of this finding is that mechanisms expressible in C-RASP may be more robustly learnable for transformers. We thus hypothesize that it is better to pre-pretrain on formal languages that can be defined in C-RASP so that the learned representations can transfer reliably.
@@ 3 Methods
Context-free and context-sensitive languages serve as formal models of the hierarchical structure of natural language. Similarly, circuit complexity classes like FO(M) and restricted programming languages like C-RASP provide theoretical upper and lower bounds on transformer capabilities. We explore the intersection of these two hierarchies to identify optimal pre-pretraining languages. In this section, we discuss how we define pre-pretraining (§3.1) and how we create formal language pre-pretraining data (§3.2).
@@ 3.1 Defining Pre-pretraining
Formally, we train a language model using an optimizer $\mathcal{A}(\mathcal{D},t,\theta_{\text{init}})$ which returns parameters $\theta_{t}$ after $t$ timesteps. We apply $\mathcal{A}$ sequentially:
@@
Our objective is to minimize the expected loss on the pretraining dataset: $\arg\min_{\theta_{t_{1}}}\;\;\mathbb{E}[\ell(\mathcal{D}_{\text{pt}},\theta_{t_{1}})]$.
We hold $\mathcal{A}$’s hyperparameters, $t_{1}$, and $\mathcal{D}_{\text{pt}}$ fixed, and we transfer model parameters directly from pre-pretraining to pretraining. So to minimize $\ell(\mathcal{D}_{\text{pt}},\theta_{t_{1}})$, we can only change the pre-pretraining dataset $\mathcal{D}_{\text{ppt}}$ and duration $t_{0}$. We compare pre-pretraining on our proposed $\mathcal{D}_{\text{ppt}}$ datasets (§3.2) against several baselines:
  • No pre-pretraining $(t_{0}=0)$
  • Random binary strings
  • Random strings of $k$ integers
  • Held-out natural language data from the same distribution as $\mathcal{D}_{\text{pt}}$
Aside from no pre-pretraining, we pre-pretrained the baselines for $t_{0}=500$ steps, for a clear comparison against the $k$-Shuffle Dyck results (see §4). The held-out natural language baseline is different from training on $\mathcal{D}_{\text{pt}}$ for longer, because we use learning rate warmup in both pre-pretraining and pretraining.
Lower validation loss compared to the no pre-pretraining baseline would indicate that pre-pretraining on formal languages is beneficial. The random baselines help establish whether this effect is specific to the particular formal languages we study. Finally, outperforming pre-pretraining on $\mathcal{D}_{\text{pt}}$ would suggest that formal languages provide better inductive biases than the pretraining data itself.
@@
In addition to validation loss, we evaluate models on BLiMP and verbatim retrieval. BLiMP compares the likelihood assigned by the model to minimal pairs of sentences that differ only in their grammaticality (e.g., Only Bill would ever complain is grammatical, but Even Bill would ever complain is not). Accuracy is measured as the proportion of examples where the grammatical sentence is assigned higher likelihood than the ungrammatical one Warstadt et al. (2020a). Verbatim retrieval tests language modeling on text passages with repeated lists Armeni et al. (2022, 2024); the model is expected to assign a very high likelihood to the words in the second repetition of the list, such that lower loss indicates better performance. Both tasks assess models’ ability to learn and apply consistent patterns—a capability that structured formal languages might strengthen. For examples of these tasks, see Tables 3 and 4 in the Appendix.
@@
In the regime with plentiful pretraining data, an ideal pre-pretraining language should minimize the number of pre-pretraining steps $t_{0}$ required: if a formal language requires very large $t_{0}$, then only pretraining on natural language would be more practical. To help quantify efficiency, we calculate the marginal rate of substitution (MRS) between formal and natural language at 10,000 steps of natural language pretraining. In other words, if we train on 500 steps of the formal language, how many more steps does it take for the natural language-only baseline to catch up?
Formally, let $x$ be the number of pre-pretraining steps and $y$ be the number of pretraining steps, and suppose the following two pairs $(x,y)$ of training steps achieve the same final loss: $(0,~{}10,000)$ and $(500,~{}6,000)$. Then the marginal rate of substitution is
@@
Similarly, the gain in token efficiency would be $1-\frac{6,000+500}{1,0000}=35\%$. For a visualization, see Figure 8.
In our setting, a good pre-pretraining language would (1) minimize the amount of pre-pretraining steps $t_{0}$ (efficiency), and (2) increase the evaluation performance of the language model. We discuss efficiency for the low-resource setting in §9.

### v2

@@ Abstract
Pretraining language models on formal language can improve their acquisition of natural language. Which features of the formal language impart an inductive bias that leads to effective transfer? Drawing on insights from linguistics and complexity theory, we hypothesize that effective transfer occurs when two conditions are met: the formal language should capture the dependency structures present in natural language, and it should remain within the computational limitations of the model architecture. We experiment with pre-pretraining (training on formal language before natural languages) on transformers and find that formal languages capturing hierarchical dependencies indeed enable language models to achieve lower loss on natural language and better linguistic generalization compared to other formal languages. We also find modest support for the hypothesis that the formal language should fall within the computational limitations of the architecture. Strikingly, pre-pretraining reduces loss more efficiently than training on a matched amount of natural language. For a 1B-parameter language model trained on roughly 1.6B tokens of natural language, pre-pretraining achieves the same loss and better linguistic generalization with a 33% smaller token budget. Finally, we also give mechanistic evidence of transfer from formal to natural language: attention heads acquired during pre-pretraining remain crucial for the model’s performance on syntactic evaluations. [^1] Code is available at https://github.com/michahu/pre-pretraining.
@@ 1 Introduction
Language models have achieved impressive performance on many tasks, but they remain data-hungry, requiring five to six orders of magnitude more data than humans to achieve human-level performance Warstadt et al. (2023); Paul (2017). This high data requirement presents challenges for training models in low-resource settings Zhong et al. (2024); Hettiarachchi et al. (2025), understanding how language models can serve as cognitive models of language acquisition with human-like data constraints Wilcox et al. (2025), and continuing to improve models even after most of the existing natural language data has been used for pretraining Villalobos et al. (2024). Thus, data efficiency during training is an important frontier for language models.
@@
A recently-explored approach for increasing data efficiency teaches models useful inductive biases by first training them on formal languages before training on natural language Papadimitriou and Jurafsky (2020); Chiang and Lee (2022); McCoy and Griffiths (2025). We refer to this paradigm as pre-pretraining. What features of formal languages make transfer to natural language effective? Papadimitriou and Jurafsky (2023) show that within the Chomsky hierarchy, context-sensitive languages transfer best to natural language compared to simpler classes of languages. We expand on their investigation and explore an additional factor: the computational limitations of the language model’s architecture. In particular, transformers—the architecture that underlies most popular language models—cannot learn all context-sensitive languages, both in theory and practice Strobl et al. (2024); Merrill and Sabharwal (2023). In fact, within all levels of the Chomsky hierarchy, some languages are harder for transformers to learn than others, and many are impossible for them to learn Merrill et al. (2023, 2024). Can a formal language give rise to positive transfer even when it cannot be fully learned by a transformer?
In this work, we hypothesize that optimal transfer from formal to natural language in transformer language models occurs at the intersection of two theoretical hierarchies: the Chomsky hierarchy of formal languages and the circuit complexity hierarchy that bounds transformer computational power (see §3). Specifically, we hypothesize that effective pre-pretraining languages should be:
@@
    learnable by transformers in a way that generalizes to longer strings than observed in training.
To satisfy the second condition, we define our formal languages in C-RASP Yang and Chiang (2024), a restricted programming language whose functions allow transformers to exhibit length generalization Huang et al. (2025).
Our empirical results support the first part of the hypothesis and provide some support for the second part (§4). Pre-pretraining on languages with hierarchical dependencies outperforms pre-pretraining on any of the other formal languages that we tested—in fact, it outperforms pre-pretraining on a matched amount of natural language. Of the formal languages with hierarchical dependencies, those that are definable in C-RASP generally achieve equal or better performance, but they are only clearly superior on some of the tasks we evaluated.
Next, we show that when positive transfer occurs, the model reuses attention heads it learned during pre-pretraining, suggesting that mechanisms from pre-pretraining transfer to natural language (§5). Finally, we scale up our experiments to a 1B-parameter language model, and show that in pre-pretraining is effective in that size as well, increasing token efficiency by 33% (§6). Overall, we conclude that formal language pre-pretraining is an effective way to improve generalization and data efficiency, and propose a hypothesis for the particular formal languages that are most promising for this purpose.
@@ 2.1 The Chomsky Hierarchy
The Chomsky hierarchy Chomsky (1959) is a nested classification of increasingly-complex formal languages. This classification is based on the kinds of computations needed to process formal structures resembling those found in human language. For example, regular languages, the least complex, can be recognized by finite-state automata. While regular languages can capture most phenomena in natural language phonology and morphology, they are insufficient for syntax: representing the hierarchical structure of natural language syntax with a finite-state automaton would require infinitely many states Chomsky (1956). Subsequent works showed that modeling some syntactic phenomena requires not only context-free but also context-sensitive grammars Shieber (1985), though the prevalence of such phenomena may be limited.
@@
A classic context-free language is $k$-Dyck: the language of well-balanced parentheses with $k$ bracket types. For example, ([])[] is a valid $2$-Dyck string, where rounded and square parentheses are the two bracket types. $k$-Dyck is often taken as a canonical example of context-free hierarchical structure because any context-free language can be reduced to Dyck via a single transformation (inverse homomorphism) and intersection with a regular language (Chomsky and Schützenberger, 1959).
@@
Removing the constraint that Dyck braces must be well-nested, but maintaining the constraint that every opening brace must be closed and vice versa, yields $k$-Shuffle Dyck, [^2] Despite what its name might suggest, $k$-Shuffle Dyck does not randomly shuffle strings from $k$-Dyck. Every opening brace in $k$-Shuffle Dyck must still be closed by a matching closing brace later in the string; this constraint would not in general be satisfied by randomly shuffled $k$-Dyck strings. Instead, $k$-Shuffle Dyck can be defined by interleaving $k$ 1-Dyck strings with different braces (Suzgun et al., 2019), as if by riffle shuffling. We use the terminology “Shuffle Dyck” for consistency with prior work. a minimal relaxation of $k$-Dyck that is strictly context-sensitive rather than context-free Suzgun et al. (2019); Strobl et al. (2024). Crossing braces in $k$-Shuffle Dyck can be thought of as a formal model of the cross-serial dependencies underlying aspects of language argued to be context-sensitive (Papadimitriou and Jurafsky, 2023).
@@ 2.2 The Circuit Hierarchy
We focus in this work on transformer language models. There are languages at each level of the Chomsky hierarchy that a transformer cannot recognize (Merrill and Sabharwal, 2023; Liu et al., 2024; Strobl et al., 2024). Thus, the Chomsky hierarchy alone does not precisely capture how difficult a language is for transformers to learn: for instance, transformers can learn some context-free languages Butoi et al. (2025) and yet fail to learn other regular languages Merrill et al. (2024). To better understand the expressive power of transformers, recent work has analyzed formal languages within a different hierarchy: the circuit complexity hierarchy, which better captures the computations performed by transformers Hao et al. (2022); Yang et al. (2024). Here, we will focus on two logics that emerge from the circuit complexity viewpoint: FO(M) Merrill and Sabharwal (2023) and C-RASP Yang and Chiang (2024).
@@
First-order logic with majority, or FO(M), is a provable upper bound on the languages that transformers can express: that is, any transformer that recognizes a language can be converted into an FO(M) program that defines (or recognizes) the same language (Merrill and Sabharwal, 2023). FO(M) programs operate by computing counts over the number of indices in an input string that satisfy certain predicates. For example, $Q_{a}(i)$ is a basic predicate that checks whether input token $i$ is an $a$. The following FO(M) program uses $Q_{a}(i)$ to define the language of strings with exactly 3 $a$’s:
@@
Beyond this example, FO(M) can implement a rich variety of programs by nesting quantifiers and building complex predicates out of logical (,,$\wedge$,$\vee$,$\neg$ , ∨ , ¬) and arithmetic operators (+,=,<+,=,<+ , = , <). In particular, FO(M) can define the $k$-Dyck language for any $k\geq 1$ (Proposition A.6 in the Appendix). For example, the following program defines $1$-Dyck:
@@
To finish the definition, we add a condition to enforce that any open and close brace paired by $\mathsf{depth}$ and $\mathsf{dindex}$ also match in their type (i.e., both are parentheses or both are square braces). See Proposition A.7 for further details.
@@
While any transformer can be compiled into FO(M), it is not necessarily the case that any FO(M) program can be implemented by a transformer. C-RASP is a restriction of FO(M) designed to be a lower bound on what transformers can express: that is, if a language is definable in C-RASP, then there exists a transformer that recognizes it (Yang and Chiang, 2024). [^3] C-RASP is a well-defined variant of the Restricted Access Sequence Processing programming language (RASP; Weiss et al., 2021; Lindner et al., 2023). The most crucial restriction for our purposes is that each C-RASP predicate can only refer to one index variable $i$, whereas in FO(M) predicates can refer to two (or more) indices $i,j$ introduced by different quantifiers (for more detail, see Yang and Chiang, 2024). This means C-RASP can define (1) or (2) above, but not $k$-Dyck for $k\geq 2$, as C-RASP cannot express the function $\mathsf{dindex}$ in (3), which compares the depth of two different indices.
Recent work has also suggested a connection between C-RASP and transformers’ ability to generalize to strings longer than those observed in training (Zhou et al., 2024; Huang et al., 2025): the definability of a language $L$ in C-RASP predicts whether transformers can reliably length-generalize when trained on strings from $L$. One interpretation of this finding is that mechanisms expressible in C-RASP may be more robustly learnable by transformers. We thus hypothesize that we will observe more reliable transfer from pre-pretraining transformers on formal languages that can be defined in C-RASP compared to languages that cannot.
@@ 3.1 Defining Pre-pretraining
We train a language model using an optimizer $\mathcal{A}(\mathcal{D},t,\theta_{\text{init}})$ which returns parameters $\theta_{t}$ after $t$ timesteps (gradient updates). We apply $\mathcal{A}$ sequentially:
@@
Our objective is to minimize the expected loss on the pretraining dataset, i.e. to find $\arg\min_{\theta_{t_{1}}}\;\;\mathbb{E}[\ell(\mathcal{D}_{\text{pt}},\theta_{t_{1}})]$. We hold $\mathcal{A}$’s hyperparameters, $t_{1}$, and $\mathcal{D}_{\text{pt}}$ fixed, and we transfer model parameters directly from pre-pretraining to pretraining. In other words, to minimize $\ell(\mathcal{D}_{\text{pt}},\theta_{t_{1}})$, we can only change the pre-pretraining dataset $\mathcal{D}_{\text{ppt}}$ and duration $t_{0}$. We compare pre-pretraining on our proposed $\mathcal{D}_{\text{ppt}}$ datasets (§3.2) against several baselines:
  • No pre-pretraining $(t_{0}=0)$.
  • Pre-pretraining on random binary strings.
  • Pre-pretraining on random strings of $k$ integers.
  • Pre-pretraining on unseen natural language data $\mathcal{D}_{\text{pt}}^{*}$ drawn from the same distribution as $\mathcal{D}_{\text{pt}}$.
Aside from the no-pre-pretraining baseline, we pre-pretrained the baselines for $t_{0}=500$ steps, the optimal number of steps for $k$-Shuffle Dyck (see §4). We note that the natural language pre-pretraining baseline is not equivalent to training on the union of $\mathcal{D}_{\text{pt}}^{*}$ and $\mathcal{D}_{\text{pt}}$ for longer, since pre-pretraining on natural language uses learning rate warmup twice, once in pre-pretraining and once in pretraining.
Lower validation loss compared to the no-pre-pretraining baseline would indicate that pre-pretraining on formal languages is beneficial. The random string baselines help establish whether this effect is specific to the particular formal languages we study. Finally, outperforming pre-pretraining on $\mathcal{D}_{\text{pt}}^{*}$ would suggest that formal languages provide a better initialization for pretraining than the pretraining data itself.
@@
In addition to measuring validation loss, we perform targeted evaluations for grammaticality and verbatim retrieval. For grammaticality judgments, we compare the likelihood assigned by the model to minimal pairs of sentences that differ only in their grammaticality (e.g., Only Bill would ever complain is grammatical, but Even Bill would ever complain is not). Accuracy is measured as the proportion of pairs where the grammatical sentence is assigned higher likelihood than the ungrammatical one Marvin and Linzen (2018). We use the BLiMP grammaticality judgment dataset Warstadt et al. (2020a). Verbatim retrieval tests language modeling on text passages with repeated lists Armeni et al. (2022, 2024); the model is expected to assign a very high likelihood to the words in the second repetition of the list, such that lower loss indicates better performance. Both evaluations assess models’ ability to learn and apply consistent patterns—a capability that could benefit from pre-pretraining on formal languages might strengthen. For examples of these evaluations, see Tables 3 and 4 in the Appendix.
@@
In the regime with plentiful pretraining data, an ideal pre-pretraining language should minimize the number of pre-pretraining steps $t_{0}$ required: if a formal language requires very large $t_{0}$ for effective transfer, then simply pretraining on natural language, without any pre-pretraining, would be more practical in terms of total compute (though even in this case pre-pretraining may still be beneficial when the amount of data available for pretraining is small, for example in low-resource languages). We quantify efficiency using the marginal rate of substitution (MRS) between formal and natural language at 10,000 steps of natural language pretraining. In other words, we ask: if we train on 500 steps of the formal language, how many more steps does it take for the natural-language-only baseline to catch up?
For example, let $x$ be the number of pre-pretraining steps and $y$ be the number of pretraining steps, and suppose the following two pairs $(x,y)$ of training steps achieve the same final loss: $(0,~{}10,000)$ and $(500,~{}6,000)$. Then the marginal rate of substitution is
@@
The gain in token efficiency would be
$1-\frac{6,000+500}{1,0000}=35\%$
For a visualization, see Figure 7.
In our setting, a good pre-pretraining language would (1) minimize the amount of pre-pretraining steps $t_{0}$ (efficiency), and (2) increase the evaluation performance of the language model.